the average of five numbers is 25. When a fifth one is added to the mix the average is 24. What is the value of the fifth number?
Answer by Alan3354(69443)

the average of five numbers is 25. When a fifth one is added to the mix the average is 24. What is the value of the fifth number?
-------
5*25 = 125
5*25 + n = 6*24 = 144
n = 19